# Accuracy of the Ensemble Kalman Filter in the Near-Linear Setting

### v1

@@ B.3 Stability Results
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{1/2}+\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}.$ (D.3)
The proof of [le2009large, Theorem 5.2] uses the law of large numbers in $L^{2}$ to establish the bound for the second term of the right hand side of D.3
$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{J}^{\rm EK}\bigr{)},R_{\phi}\Bigr{)}}{\sqrt{N}};$ (D.4)
furthermore, in the proof of [le2009large, Theorem 5.2], the following bound on the first term of the right hand side of D.3 may be deduced:
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{1/2}\leqslant L_{\phi}D_{J}^{N,4}\cdot\Bigl{(}1+\bigl{|}\mathcal{M}^{1/4}_{4\varsigma}\bigl{(}\mu_{J}^{\rm EK}\bigr{)}\bigr{|}\Bigr{)}+L_{\phi}|D_{J}^{N,2(\varsigma+1)}|^{\varsigma+1}.$ (D.5)
Under the assumptions of the proposition, the result of [le2009large, Proposition 4.4] shows by induction that for any $j\in\mathbb{N}$, it holds that  [^1] The result of [le2009large, Proposition 4.4] holds more generally for functions $\Psi$ that are locally Lipschitz and that grow at most polynomially at infinity [le2009large, Assumption B]. However, to obtain (D.6) from the result of [le2009large, Proposition 4.4] we use that $\Psi$ is assumed to be globally Lipschitz.
$D_{j}^{N,p}\leqslant CD_{j-1}^{N,p}$ $+CD_{j-1}^{N,2\vee(pr)}\Bigl{(}D_{j-1}^{N,2\vee(pr^{\prime})}+\bigl{(}\mathcal{M}\bigl{(}\mathsf{P}\mu_{j-1}^{\rm EK}\bigr{)}\bigr{)}^{2\vee(pr^{\prime})}\Bigr{)}\cdot\Bigl{(}D_{j-1}^{N,2\vee(pr^{\prime\prime})}+\overline{S}_{j}^{pr^{\prime\prime}}\Bigr{)}$ (D.6)
$\quad+C\Bigl{(}\mathbb{E}\bigl{|}Z_{j}^{N}\bigr{|}^{pq}\Bigr{)}^{1/pq}\cdot\bigl{(}D_{j-1}^{N,pq^{\prime}}+\overline{S}_{j}^{pq^{\prime}}\bigr{)},$
for any $j\in\mathbb{N}$, where $C$ is a constant depending on $\Sigma,\Gamma,\ell_{\Psi},\kappa_{\Psi},\kappa_{h},\mathcal{M}_{2}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)}$. We note that the specific form of $C$ may be deduced using the proof of [le2009large, Lemma 4.1]. Furthermore, we note that
$\frac{1}{q}+\frac{1}{q^{\prime}}=1,\qquad\frac{1}{r}+\frac{1}{r^{\prime}}+\frac{1}{r^{\prime\prime}}=1,$
and that $\overline{S}_{j}^{t}$ and $Z_{j}^{N}$ are defined as
$\overline{S}_{j}^{t}$ $\coloneqq\Bigl{(}\mathbb{E}\bigl{|}{y}^{\dagger}_{j}-h\Bigl{(}\overline{v}^{(i)}_{j}\Bigr{)}-\eta^{(i)}_{j}\bigr{|}^{t}\Bigr{)}^{1/t},$
$Z_{j}^{N}$ $\coloneqq\left\|\mathcal{C}\Bigl{(}\frac{1}{N}\sum_{i=1}^{N}\delta_{\overline{v}^{(i)}_{j}}\Bigr{)}-\mathcal{C}\bigl{(}\mathsf{P}\mu_{j-1}^{\rm EK}\bigr{)}\right\|.$
By applying the triangle inequality it is straightforward to deduce that
$\overline{S}_{j}^{t}\leqslant\left(\int_{{}^{d_{u}}}\bigl{|}{y}^{\dagger}_{j}-h(u)\bigr{|}^{t}\mathsf{P}\mu_{j-1}^{\rm EK}(\mathrm{d}u)\right)^{1/t}+\bigl{(}\mathbb{E}|\eta_{j}|^{t}\bigr{)}^{1/t}\leqslant C\Bigl{(}\mathcal{M}_{t}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)},\kappa_{\Psi},\kappa_{h},\kappa_{y},\Sigma,\Gamma\Bigr{)},$ (D.7)
for any $t\in\mathbb{N}$. On the other hand, [le2009large, Lemma 3.3] shows that
$\Bigl{(}\mathbb{E}\bigl{|}Z_{j}^{N}\bigr{|}^{t}\Bigr{)}^{1/t}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2t}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)},\kappa_{\Psi},\Sigma\Bigr{)}}{\sqrt{N}},$ (D.8)
for any $N\in\mathbb{N}$ and any $t\in\mathbb{N}$. Combining the results from (D.7) and (D.8) with the one in (D.6), and choosing for convenience $q=q^{\prime}=2$ and $r=r^{\prime}=r^{\prime\prime}=3$ for any $j\in\mathbb{N}$ we obtain
$D_{j}^{N,p}\leqslant CD_{j-1}^{N,p}+CD_{j-1}^{N,3p}\Bigl{(}D_{j-1}^{N,3p}+1\Bigr{)}\cdot\Bigl{(}D_{j-1}^{N,3p}+1\Bigr{)}+\frac{C}{\sqrt{N}}\cdot\bigl{(}D_{j-1}^{N,2p}+1\bigr{)},$ (D.9)
for a constant $C$ depending on $\mathcal{M}_{4p}{\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)}},\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$. Since $D_{0}^{N,p}=0$ for any $p\in\mathbb{N}$, by iterating in $j$ the result in (D.9), we find that
$D_{J}^{N,4}\leqslant\frac{C}{\sqrt{N}},$
where $C$ is a constant depending on $\mathcal{M}_{16\cdot 3^{J-1}}{\bigl{(}\mu_{0}^{\rm EK}\bigr{)}},J,\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$. Hence, by combining the results of (D.5) and (D.4) with the decomposition (D.3), we deduce that
@@
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}},$
@@
where $C$ is a constant depending on $\mathcal{M}_{q}{\bigl{(}\mu_{0}^{\rm EK}\bigr{)}},J,R_{\phi},L_{\phi},\varsigma,\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$ for a constant $q$ defined as $q\coloneqq\max\{16\cdot 3^{J-1},8(\varsigma+1)3^{J-1}\}$. ∎
[…]

### v2

@@ B.3 Stability Results
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}$ $\leqslant\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$
@@
$\qquad+\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}.$ (D.3)
Step 1: bounding the second term in (D.3)
Noting that the random variables $\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu_{J}^{{\rm EK}}[\phi]$ are i.i.d. and expanding the square, we obtain
$\left(\mathbb{E}~{}\biggl{\lvert}\sum_{i=1}^{N}\frac{1}{N}\Bigl{(}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{{\rm EK}}_{J}[\phi]\Bigr{)}\biggr{\rvert}^{2}\right)^{\frac{1}{2}}=\frac{1}{\sqrt{N}}\Bigl{(}\mathbb{E}\bigl{|}\phi\bigl{(}\overline{u}^{(1)}_{J}\bigr{)}-\mu_{J}^{{\rm EK}}[\phi]\bigr{|}^{2}\Bigr{)}^{\frac{1}{2}},$
Since by Assumption P1 it holds that $|\phi(u)|\leqslant R_{\phi}\bigl{(}1+|u|^{{\varsigma}+1}\bigr{)}$ for any $u\in^{d_{u}}$, it follows that
$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},R_{\phi},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}}{\sqrt{N}},$ (D.5)
where we used Lemma B.5 to bound $\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{J}^{{\rm EK}}\bigr{)}$ in terms of $\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)}$.
Step 2: bounding the first term in (D.3)
For the first term, by Jensen’s inequality, exchangeability, and finally by Assumption P1 on $\phi$ together with the Cauchy–Schwarz inequality, it holds without loss of generality that
$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\Bigl{(}u^{(i)}_{J}\Bigr{)}-\phi\Bigl{(}\overline{u}^{(i)}_{J}\Bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$ $\leqslant\left(\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\Bigl{|}\phi\Bigl{(}u^{(i)}_{J}\Bigr{)}-\phi\Bigl{(}\overline{u}^{(i)}_{J}\Bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$
$=\left(\mathbb{E}\Bigl{|}\phi\Bigl{(}u^{(1)}_{J}\Bigr{)}-\phi\Bigl{(}\overline{u}^{(1)}_{J}\Bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$
$\leqslant 3L_{\phi}\left(\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{4}\right)^{\frac{1}{4}}\left(\mathbb{E}\left[1+\left\lvert u^{(1)}_{J}\right\rvert^{4\varsigma}+\left\lvert\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\right]\right)^{\frac{1}{4}}.$ (D.6)
@@
By Lemma B.5, there exists $C=C\Bigl{(}\mathcal{M}_{4\varsigma}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ so that
$\mathbb{E}\left\lvert\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\leqslant C\quad\text{and}\quad\mathbb{E}\left\lvert u^{(1)}_{J}\right\rvert^{4\varsigma}\leqslant 2^{4\varsigma-1}\left(\mathbb{E}\left\lvert\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}+\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\right)\leqslant 2^{4\varsigma-1}\left(C+\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\right).$
Hence it remains to bound terms of the form $\mathbb{E}\bigl{\lvert}u^{(1)}_{J}-\overline{u}^{(1)}_{J}\bigr{\rvert}^{p}$ with $p=4$ and $p=4\varsigma$. Such a bound will follow from a propagation of chaos result; which will be the object of Step 4. In Step 3 we show auxiliary moment bounds for the mean field dynamics.
Step 3: moment bounds for the mean field dynamics
By Lemma B.5, for any $j\in\llbracket 0,J-1\rrbracket$ there exists a constant $C=C\Bigl{(}\mathcal{M}_{p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ so that
$\left(\mathbb{E}\left\lvert\overline{u}_{j}^{(1)}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C.$ (D.7)
From this it immediately follows, by Assumptions (H3) and (H4), that there exist constants $C_{v},C_{y}$ depending on $\mathcal{M}_{p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$ so that
$\left(\mathbb{E}\left\lvert\overline{v}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{v},\qquad\left(\mathbb{E}\left\lvert\overline{y}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{y}.$ (D.8)
Furthermore, it holds that for the Kalman gain that
$\Bigl{\lVert}\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\Bigr{\rVert}^{p}\leqslant\left\lVert\mathcal{C}^{uh}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\right\rVert^{p}\left\lVert\Gamma^{-1}\right\rVert^{p}=\left\lVert\Gamma^{-1}\right\rVert^{p}\left\lVert\mathcal{C}^{uu}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}H^{\mathsf{T}}\right\rVert^{p}\leqslant C,$ (D.9)
where $C$ depends on $\mathcal{M}_{2}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
Step 4: propagation of chaos
In this step, we prove a propagation of chaos result stating for all $p$, there exists a constant $C_{p}=C_{p}\Bigl{(}\mathcal{M}_{4^{J}\cdot p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma\Bigr{)}$ such that
$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{p}.$ (D.10)
To prove this result, we follow the strategy in [le2009large] and reason by induction. Fix $j\in\llbracket 0,J-1\rrbracket$ and assume that for all $p\in\mathbb{N}$ there is $C_{j,p}=C_{j,p}\Bigl{(}\mathcal{M}_{4^{j}\cdot p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma\Bigr{)}$ such that
$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert u^{(1)}_{j}-\overline{u}^{(1)}_{j}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{j,p}.$ (D.11)
By Lipschitz continuity of $\Psi$ and $h$, we deduce immediately that
$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert v^{(1)}_{j+1}-\overline{v}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant\ell_{\Psi}C_{j,p},\qquad\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert y^{(1)}_{j+1}-\overline{y}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant\ell_{h}\ell_{\Psi}C_{j,p}.$ (D.12)
Now we write
$u^{(1)}_{j+1}-\overline{u}^{(1)}_{j+1}$ $=\left(v^{(1)}_{j+1}-\overline{v}^{(1)}_{j+1}\right)+\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\left(\overline{y}^{(1)}_{j+1}-y^{(1)}_{j+1}\right)$
$\qquad+\Bigl{(}\mathcal{K}\bigl{(}\pi^{{\rm EK},N}_{j+1}\bigr{)}-\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\Bigr{)}\left(y^{\dagger}_{j+1}-y^{(1)}_{j+1}\right).$
Thus, by the triangle and Hölder inequalities,
$\sqrt{N}\left(\mathbb{E}\left\lvert u^{(1)}_{j+1}-\overline{u}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}$ $\leqslant\ell_{\Psi}C_{j,p}+\left\lVert\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\right\rVert\ell_{h}\ell_{\Psi}C_{j,p}$
$\qquad+\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\bigl{(}\pi^{{\rm EK},N}_{j+1}\bigr{)}-\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}\left(\mathbb{E}\left\lvert y^{\dagger}_{j+1}-y^{(1)}_{j+1}\right\rvert^{2p}\right)^{\frac{1}{2p}}.$
The first two terms on the right-hand side are bounded uniformly in $N$ in view of (D.9). In order to bound the last term, we note that the term $\mathbb{E}\bigl{\lvert}y^{\dagger}_{j+1}-y^{(1)}_{j+1}\bigr{\rvert}^{2p}$ is bounded uniformly in $N$ by a constant which depends on $\mathcal{M}_{4^{j}\cdot 2p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)}$; this follows by (D.8) and (D.12). To complete the inductive step, it remains to show that
$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}<\infty.$
To this end, in line with the classical propagation of chaos approach, we decompose
$\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}$ $\leqslant\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}$
$\qquad+\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}},$ (D.13)
where we introduced the empirical measure associated with the mean field particles:
$\overline{\pi}^{{\rm EK},N}_{j+1}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{\left(\overline{v}^{(i)}_{j+1},\overline{y}^{(i)}_{j+1}\right)}.$
To conclude the proof of propagation of chaos, we must prove that
  • The first term on the right-hand side of (D.13) is bounded uniformly in $N$, which is achieved by employing the induction hypothesis (D.11), as well as (D.12).
  • The second term on the right-hand side of (D.13) is also bounded uniformly in $N$. This follows from classical arguments based on the law of large number in $L^{p}$ spaces.
Let us first bound the second term. To this end, note that
$\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}$ $=\operatorname{\mathcal{C}}^{uh}\left(\overline{\pi}^{{\rm EK}}_{j+1}\right)\left(\left(\Gamma+\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right)^{-1}-\left(\Gamma+\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right)^{-1}\right)$
$\qquad+\left(\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right)\left(\Gamma+\operatorname{\mathcal{C}}^{hh}\left(\overline{\pi}^{{\rm EK},N}_{j+1}\right)\right)^{-1}.$
Therefore, using that $A^{-1}-B^{-1}=B^{-1}(B-A)A^{-1}$, we have that
$\left\lVert\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert$ $\leqslant\left\lVert\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert\lVert\Gamma^{-1}\rVert^{2}\left\lVert\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert$
$\qquad+\lVert\Gamma^{-1}\rVert\left\lVert\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert$
$\leqslant C(\Gamma,H)\Bigl{(}1+\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert\Bigr{)}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert.$
The term $\bigl{\lVert}\operatorname{\mathcal{C}}^{uu}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\bigr{\rVert}$ is bounded by (D.8), and the $L^{p}$ norm of the second term tends to zero with rate $N^{-\frac{p}{2}}$ by classical law of large number arguments, see for example [MR4234152, Lemma 3] and [vaes2024, Lemma 3]; indeed it holds that
$\mathbb{E}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\leqslant\frac{C\Bigl{(}\mathcal{M}_{4p}\bigl{(}\mu^{{\rm EK}}_{0}\bigr{)}\Bigr{)}}{N^{p}}.$
It remains to bound the other term. Reasoning similarly to above, we have that
$\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert\leqslant C(\Gamma,H)\Bigl{(}1+\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert\Bigr{)}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert.$ (D.14)
By [vaes2024, Lemma 2], it holds that
$\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert^{2}\leqslant 2\left(\frac{1}{N}\sum_{i=1}^{N}\left\lvert v^{(i)}_{j+1}-\overline{v}^{(i)}_{j+1}\right\rvert^{2}\right)\left(\frac{1}{N}\sum_{i=1}^{N}\left\lvert v^{(i)}_{j+1}\right\rvert^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\lvert\overline{v}^{(i)}_{j+1}\right\rvert^{2}\right).$
Using the Cauchy-Schwarz and Jensen’s inequalities, moment bounds and finally the induction hypothesis (D.10) enables to conclude that
$\left(\mathbb{E}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}\leqslant\frac{C\Bigl{(}\mathcal{M}_{4^{j+1}p}\bigl{(}\mu^{{\rm EK}}_{0}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma\Bigr{)}}{\sqrt{N}}.$
Step 5: concluding the proof
Using the propagation of chaos result from (D.10) in (D.6) gives that
$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\left(u^{(i)}_{J}\right)-\phi\left(\overline{u}^{(i)}_{J}\right)\Bigr{|}^{2}\right)^{\frac{1}{2}}\leqslant\frac{C}{\sqrt{N}},$
for some constant $C$ depending on $\mathcal{M}_{q}\bigl{(}\mu^{{\rm EK}}_{0}\bigr{)},J,L_{\phi},\varsigma,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma$ with $q=\max\{4^{J+1},4\varsigma\}$. Combining this result with (D.5) yields the desired result. ∎